Circular No. 7078

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE NOVA IN MUSCA W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag 8.5) on Technical Pan photographs taken with PROBLICOM on Dec. 29.3090 and 29.3123 UT with an 85-mm camera lens (+ orange filter); he gave the location of the new star as α = 11^{h}56^{m}.2, δ = −65^{o}34′ (equinox 2000.0). Nothing brighter than mag 11.5 appears at this location on photographs taken on Dec. 20. A CCD observation by Liller shows the star was at mag V = 9.45 on Dec. 30.1465. A. Pearce, Nedlands, W. Australia, estimated m_{v} = 10.0 on Dec. 30.80, providing the position as α = 11^{h}56^{m}.2, δ = −65^{o}33′.

GRB 981226 F. Frontera and E. Montanari, Universitá, Ferrara; L. A. Antonelli,

Osservatorio Astronomico, Rome; P. Giommi, Agenzia Spaziale Italiana, Rome; L. Piro, IAS, CNR, Rome; M. R. Daniele, A. Coletta, and C. De Libero, BeppoSAX, Rome, on behalf of a large collaboration, report: “The BeppoSAX Wide Field Camera error box of GRB 981226 (IAUC 7074) was observed with the BeppoSAX Narrow Field Instruments (NFI) starting ∼ 11 hr after the burst in the period Dec. 26.878–29.148 UT. Preliminary analysis of MECS2+3 data shows a previously unknown x-ray source, 1SAX J2329.6–2356, located at α = 23^{h}29^{m}36^{s}.0, δ = −23^{o}55′58′′ (equinox 2000.0), with an error radius of 1′. During the first 33.3 hr of observation, the 2–10-keV source flux decreases by a factor 3, from 0.0043 ± 0.0010 counts/s, corresponding to (3.0 ± 0.7) × 10^{−13} erg^{−1} cm^{−2}, down to 0.00167 ± 0.00046 counts/s, corresponding to (0.98 ± 0.30) × 10^{−13} erg^{−1} cm^{−2}. The measurements during the subsequent 18.36 hr show that the flux of 1SAX J2329.6–2356 does not vary significantly from the last value. Very likely this source is the x-ray afterglow of GRB 981226.”

COMET P/1998 Y2 (LI) As indicated on IAUC 7075, additional observations have shown that

this is indeed a short-period comet with P ∼ 15.2 yr. The revised orbit has been issued on MPEC 1998-Y41.

SUPERNOVAE 1998es AND 1998ey CCD R magnitudes by E. Prosperi, Larciano, Italy: 1998es in NGC

632, Dec. 26.84 UT, 14.6; 1998ey in NGC 7080, Dec. 26.73, 16.8.

1998 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green